#4 Assume a par value of $1,000. Caspian Sea plans to issue a 10.00 year, semi-annual...

#4
Assume a par value of $1,000. Caspian Sea plans to issue a 10.00 year, semi-annual pay bond that has a coupon rate of 8.01%. If the yield to maturity for the bond is 7.71%, what will the price of the bond be?


Submit
Answer format: Currency: Round to: 2 decimal places.


unanswered
not_submitted
#5
Assume a par value of $1,000. Caspian Sea plans to issue a 11.00 year, semi-annual pay bond that has a coupon rate of 7.92%. If the yield to maturity for the bond is 8.27%, what will the price of the bond be?


Submit
Answer format: Currency: Round to: 2 decimal places.
Answer #1

Hello Sir/ Mam

#4

Assume a par value of $1,000. Caspian Sea plans to issue a 10.00 year, semi-annual pay bond that has a coupon rate of 8.01%. If the yield to maturity for the bond is 7.71%, what will the price of the bond be?

Using excel function, "=PV(7.71%/2,10*2,-80.1/2,-1000,0)", we get PV = $1,020.65

#5

Assume a par value of $1,000. Caspian Sea plans to issue a 11.00 year, semi-annual pay bond that has a coupon rate of 7.92%. If the yield to maturity for the bond is 8.27%, what will the price of the bond be?

Using excel function, "=PV(8.27%/2,11*2,-79.2/2,-1000,0)", we get PV = $975.03

I hope this solves your doubt.

Feel free to comment if you still have any query or need something else. I'll help asap.

Do give a thumbs up if you find this helpful.